Smith purchased 10 dollar stamp, 15 dollar stamps and 20 dollars with a total value of 840 dollars. If the number of 15 dollars stamps is 8 more than 10 dollar stamps and there are 10 more of the 20 dollar stamps than of the 15 dollar stamps, how many of each does she
receive?

Answer by stanbon(75887)

Smith purchased 10 dollar stamp, 15 dollar stamps and 20 dollars with a total value of 840 dollars. If the number of 15 dollars stamps is 8 more than 10 dollar stamps and there are 10 more of the 20 dollar stamps than of the 15 dollar stamps, how many of each does she
receive?
-----
Equations
10t + 15f + 20w = 840 dollars
----
f = t + 8
w = f + 10
------
Substitute for t and w and solve for "f"::
10(f-8) + 15f + 20(f+10) = 840
45f + 120 = 840
45f = 720
f = 16 (# of 15 dollar bills)
-----
t = f-8 = 8(# of ten dollar bills)
-----
w = f + 10 = 26 (# of twenty dollar bills)
---------------
Cheers,
Stan H.
---------------